# DGC-GNN: Descriptor-free Geometric-Color Graph Neural Network for 2D-3D Matching

### v1

@@ 2.2.3 Cluster-based Local Matching
After extracting the global geometric embedding, we implement a cluster-based matching module to obtain the initial intra-domain 2D-3D matches. This cluster-based GNN  shi2022clustergnn has been shown to be more computationally efficient than its complete-graph counterpart sarlin2020superglue . The network considers the local point features from both $\mathcal{\widetilde{P}}$ and $\mathcal{\widetilde{Q}}$ a complete set, then classifies the local feature with strong correlations into the same group and restricts the message passing within each group. In addition to its low computational complexity, we found that the cluster GNN can effectively utilize our global-to-local geometric cues, as the clustering operation inherits the property of global graph clustering and forces it to distinguish ambiguous local features even with similar global embedding.
@@
GoMatch-style Initialization. Our local graph initialization follows the pipeline in GoMatch zhou2022geometry and conducts it as geometric self-attention. For each local point $\mathbf{b}_{\mathbf{p},n}$, we construct a local graph according to its $k^{\prime}$ nearest neighbours in the Euclidean space and update the associated feature $\mathbf{\widetilde{f}}_{\mathbf{p},n}\in\mathbb{R}^{2d}$ by Eq. 4. Note that we ignore the angular embedding at this stage due to the unaffordable memory requirements with space complexity $\mathcal{O}(Nk^{\prime 2})$, where $N$ is the number of local points. We then adopt linear attention katharopoulos2020transformers ; sun2021loftr as a cross-attention mechanism, which allows each keypoint in one modality to interact with all keypoints from another modality. This not only facilitates inter-modality in the feature matching but also reduces the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$.
@@
Cluster-based Attention. After the graph initialization, the features $\mathbf{\widetilde{f}}_{\mathbf{p}}$ and $\mathbf{\widetilde{f}}_{\mathbf{q}}$ coming from the image and the point cloud respectively, are concatenated and processed in a two-level hierarchical clustering attention module. The hierarchical structure is effective in reducing erroneous grouping. At the first level, we directly cluster the feature vectors into $I$ coarse groups. In the second level, each coarse group is divided into several small groups. The local point information exchange is conducted at each level and only within each group to obtain more representative features. After the sparse clustering, each feature vector is transformed back to its original position and then split again into $\mathbf{\widetilde{f}^{\prime}}_{\mathbf{p}}$ and $\mathbf{\widetilde{f}^{\prime}}_{\mathbf{q}}$ to obtain the keypoints both in the 2D and 3D spaces.
@@ 2.3 Outlier Rejection
After obtaining the initial matches, an outlier pruning approach is adopted to remove the incorrect ones. We apply a classification network, as applied in yi2018learning ; zhou2022geometry , whose input is the concatenated 2D and 3D keypoint features $\mathbf{\widetilde{f}^{\prime}}_{\widetilde{n},\widetilde{m}}=\mathbf{\widetilde{f}^{\prime}}_{\mathbf{p},\widetilde{n}}\oplus\mathbf{\widetilde{f}^{\prime}}_{\mathbf{q},\widetilde{m}}$ and outputs the matching confidence of each matched pair. The final predicted matches are obtained as follows:
$\mathcal{M}_{final}=\{(\widetilde{n}^{\prime},\widetilde{m}^{\prime})\ |\ \forall\ \operatorname{classifier}(\mathbf{\widetilde{f}^{\prime}}_{\widetilde{n},\widetilde{m}}\;|\;(\widetilde{n},\widetilde{m})\in\mathcal{M}_{init})\geq\theta\},$ (8)
@@ 2.4 Training Loss
$\mathcal{L}_{ot}=-\frac{1}{|\mathcal{M}_{gt}|+|\mathcal{I}|+|\mathcal{J}|}(\sum\limits_{(n,m)\in\mathcal{M}_{gt}}\log\mathcal{\bar{S}}_{n,m}+\sum\limits_{n\in\mathcal{I}}\log\mathcal{\bar{S}}_{i,m+1}+\sum\limits_{m\in\mathcal{J}}\log\mathcal{\bar{S}}_{N+1,j}).$ (9)
@@
where $w_{i}$ is the balance weight of positive and negative samples, $y_{i}$ is the gt matching label for the $i$-th correspondences, $p_{i}$ is the predicted probability of a true match for the $i$-th correspondences. The total loss is the sum of the two terms, thus, $\mathcal{L}=\mathcal{L}_{ot}+\mathcal{L}_{or}$.
@@ 3 Experiments
In this section, we first describe the experimental setup, employed datasets, and evaluation protocols. Next, we provide detailed comparisons with the baselines on both matching and visual localization tasks. Finally, we conduct ablation studies on each designed component.
@@
Training. We train the indoor model of DGC-GNN on the ScanNet dai2017scannet dataset and the outdoor model on the MegaDepth li2018megadepth dataset. We extract up to 1024 keypoints for each training image by the SIFT detector lowe2004distinctive . Similarly as in GoMatch, we first select a subset of the point cloud by applying image retrieval approaches Arandjelovic16 ; torii201524 to obtain potential images observing the same part of the scene as the input one. During the training, we randomly sample the retrieval pairs with a visual overlap of more than 35% on MegaDepth and 65% on ScanNet to ensure enough matches on each pair. For the global geometric embedding, we cluster the 2D/3D bearing vectors into $K=10$ groups, and each cluster center point is connected to its $k=4$ nearest neighbors to build the global graph. For the local point graph, we connect each point with its 10 nearest neighbors and the cluster-based attentions are performed twice to force the intra-cluster information exchange.
@@
The model is optimized using the Adam with a fixed learning rate of 1e-3. We train DGC-GNN with one 32GB Telsa V100 GPU. The convergence of the model typically requires 25 epochs with around 20 hours for indoor scenes and 35 epochs (approx. 35 hours) for outdoor scenes.
@@
Datasets. We use ScanNet and MegaDepth for the model training and 2D-3D matching task evaluation. As a downstream application, we perform visual localization on the 7Scenes shotton2013scene and Camdridge-Landmarks kendall2015posenet datasets. MegaDepth is a popular outdoor dataset with 196 scenes captured around the world. The sparse 3D reconstructions are provided by the COLMAP schoenberger2016sfm structure-from-motion software. Following zhou2022geometry , we train our outdoor model on 99 scenes and evaluate it with 53 scenes. ScanNet is a large-scale RGB-D indoor dataset that consists of 1613 scans with over 2.5 million images. We randomly selected 105 scenes for the training and 30 scenes for the evaluation. Cambridge-Landmarks is a middle-scale outdoor dataset consisting of 6 individual scenes. The ground truth camera poses are provided by a structure-from-motion algorithm. We follow kendall2015posenet ; zhou2022geometry to evaluate our method on four of the scenes. 7scenes is a small indoor dataset with RGB-D images and camera poses provided by the depth SLAM system. We evaluate on the standard test sequences.
@@
Evaluation Protocol. For matching on ScanNet and MegaDepth, we follow zhou2022geometry and report the AUC score calculated from the reprojection errors. To calculate the reprojection errors for the 2D-3D matches in $\mathcal{M}_{final}$, we project the 3D points to the image plane using the ground truth and estimated camera poses. Then we calculate the $L_{2}$ distance between the ground truth and estimated reprojected 2D points. We use multiple thresholds, 1, 5, and 10 pixels, to evaluate the algorithms. The camera translation and rotation error quantiles at 25%, 50%, and 75% are also reported. Besides, we also evaluate the matching quality by calculating the matching precision P, which is the ratio of inlier matches after PnP-RANSAC to the number of final matches $\mathcal{M}_{final}$. For visual localization tasks, we report the median translation (in meters) and rotation (in degrees) camera pose errors.
Methods ScanNet dai2017scannet MegaDepth li2018megadepth
BPnPNet campbell2020solving 10.00 / 10.00 / 10.02 199.17 / 128.90 / 154.68 4.35 / 6.82 / 9.86 13.60 10.22 / 10.63 / 10.89 16.13 / 32.01 / 61.58 1.67 / 3.17 / 5.44 12.95
GoMatch zhou2022geometry 11.18 / 11.23 / 18.01 112.69 / 112.78 / 136.50 0.19 / 0.91 / 2.63 13.18 15.67 / 22.43 / 28.01 10.60 / 10.08 / 34.63 0.06 / 1.06 / 3.73 14.94
k=10 BPnPNet campbell2020solving 10.00 / 10.00 / 10.03 104.68 / 135.94 / 160.54 4.67 / 7.30 / 10.92 10.84 10.36 / 10.72 / 10.97 16.63 / 34.69 / 67.77 1.64 / 3.30 / 5.97 10.74
GoMatch zhou2022geometry 10.91 / 18.98 / 31.12 111.18 / 114.94 / 128.97 10.08 / 0.35 / 2.08 14.25 18.90 / 35.67 / 44.99 10.18 / 11.29 / 16.65 0.02 / 0.12 / 1.92 18.76
Table 1: 2D-3D matching. We report the reprojection error AUC scores thresholded at 1, 5, and 10 pixels; rotation (degrees) and translation error (meters) quantiles at 25, 50, and 75%; and matching precision (percentages). On MegaDepth li2018megadepth , the translation scale is unknown, so the unit of the errors comes from the reference reconstruction and is not in meters. Parameter $k$ is the number of images retrieved from the database to narrow down the search space. The best results are bold. DGC-GNN nearly doubles the AUC scores of GoMatch and reduces the pose errors to their ${\approx}33\%$.
@@
We compare with the two descriptor-free matchers GoMatch zhou2022geometry and BPnPnet campbell2020solving . During inference, we use the 3D points from the top-$k$ retrieved database images to match against the keypoints from query images. Following zhou2022geometry , we report the upper bound of the AUC score at 1, 5, and 10 pixels by the ground truth matches. We refer to these values as Oracle. We use the official code with the default setting to generate the evaluation dataset on MegaDepth li2018megadepth and rerun the GoMatch and BPnPNet with the released models on the dataset. Note that we also tested GoMatch after retraining it on MegaDepth and achieved similar results as with the released model.
Matching Results. The results with $k=1$ and $k=10$ are presented in Table 1. $k$ is the number of retrieved image pairs that are used for evaluation. The proposed method outperforms GoMatch and BPnPNet by a significant margin on both scenes. Specifically, DGC-GNN achieves 10.2 / 37.64 / 44.04% reprojection AUC compared to GoMatch with 5.67 / 22.43 / 28.01% on MegaDepth with $k=1$. DGC-GNN halves the rotation and translation errors of GoMatch on all thresholds and it obtains better matching quality. Notably, the performance of DGC-GNN with $k=1$ surpasses that of GoMatch with $k=10$, indicating the effectiveness of our method even with a single view.
Figure 4: Outlier Sensitivity. The AUC scores of BPnPNet campbell2020solving , GoMatch zhou2022geometry and the proposed DGC-GNN thresholded at 1, 5, and 10 pixels are plotted as a function of the outlier ratio. Oracle represents the AUC score upper bound using ground truth 2D-3D matches.
Sensitivity to Outliers. To evaluate the sensitivity of our method to keypoint outliers, we follow the procedure in GoMatch zhou2022geometry . The outliers are controlled by the outlier ratio, ranging from 0 to 1, calculated as the number of unmatched keypoints divided by the maximum of the numbers of 2D and 3D points. In case the outlier ratio is 00, all of the input 2D and 3D points are selected from the ground truth matches, and no outliers are included in the matching process. When it is $1$, we directly use the keypoints from the 2D query and 3D points from top-$k$ retrieved images as input to find the matches without any filtering or outlier removal. The results are shown in Fig. 4. Even in the presence of outliers, DGC-GNN outperforms other methods by a large margin. This indicates that our method is more robust to the presence of outliers and can handle challenging matching scenarios more effectively than the state-of-the-art.
@@
Table 2: Ablation Study. AUC scores thresholded at 1, 5, and 10 pixels; rotation and translation error quantiles at 25, 50, 75% with the proposed components added one by one to the GoMatch pipeline.
Ablation Study. We investigate the effectiveness of different components of DGC-GNN on the 2D-3D matching quality. The impact of different components of DGC-GNN is studied on MegaDepth dataset li2018megadepth with $k=10$. The results are reported in Table 2. We provide other results in the supplementary material. We conduct the ablations by gradually adding the components: global geometric embedding (G. Emb), cluster attention (C. Att.), Color, and Angular embedding (Ang.) to the original GoMatch pipeline. Incorporating color information into the matching process significantly impacts the performance, resulting in improvements of 2.55 / 3.88 / 2.55% (AUC@1 / 5 / 10px). This demonstrates the importance of considering color cues for accurate and robust matching. The global-to-local geometric embedding (G. Emb.) and Angular relation embedding (Ang.) substantially improve performance by 1.90 / 5.51 / 5.52% and 1.11 / 3.36 / 3.47%, respectively. It highlights the effectiveness of our idea incorporating global geometric context and local geometric details. The cluster attention mechanism also plays a vital role, improving performance by 0.78 / 3.28 / 3.48%. The best results are obtained when all components are added to the pipeline.
Methods No Desc. Maint. Privacy Cambridge-Landmarks kendall2015posenet (cm, ${}^{\circ}$) MB used 7Scenes shotton2013scene (cm, ${}^{\circ}$) MB used
@@
E2E MS-Trans. shavit2021learning 83 / 1.47 181 / 2.39 86 / 3.07 162 / 3.99 1171 11 / 4.66 24 / 9.60 14 / 12.19 17 / 5.66 18 / 4.44 17 / 5.94 26 / 8.45 11171
DSAC* brachmann2021dsacstar 15 / 0.30 121 / 0.40 15 / 0.30 113 / 0.40 1112 12 / 1.10 12 / 1.24 11 / 1.82 13 / 1.15 14 / 1.34 14 / 1.68 13 / 1.16 11196
HSCNet li2020hierarchical 18 / 0.30 119 / 0.30 16 / 0.30 119 / 0.30 1592 12 / 0.70 12 / 0.90 11 / 0.90 13 / 0.80 14 / 1.00 14 / 1.20 13 / 0.80 11036
DB HybridSC camposeco2019hybrid 81 / 0.59 175 / 1.01 19 / 0.54 150 / 0.49 1113 - - - - - - - -
AS sattler2016efficient 13 / 0.22 120 / 0.36 14 / 0.21 118 / 0.25 1813 13 / 0.87 12 / 1.01 11 / 0.82 14 / 1.15 17 / 1.69 15 / 1.72 4 / 1.01 -
SP detone2018superpoint +SG sarlin2020superglue 12 / 0.20 115 / 0.30 14 / 0.20 117 / 0.21 3215 12 / 0.85 12 / 0.94 11 / 0.75 13 / 0.92 15 / 1.30 14 / 1.40 15 / 1.47 22977
DF GoMatch zhou2022geometry 25 / 0.64 283 / 8.14 48 / 4.77 335 / 9.94 1148 14 / 1.65 13 / 3.86 19 / 5.17 11 / 2.48 16 / 3.32 13 / 2.84 89 / 21.12 11302
@@
Table 3: Visual localization. We report the median pose errors (cm, ${}^{\circ}$) and storage requirements (MB) on the scenes of the 7Scenes shotton2013scene and Cambridge-Landmarks kendall2015posenet datasets. Three groups of methods are shown: end-to-end (E2E), descriptor-based (DB), and descriptor-free (DF). We do not show BPnPNet as it fails on most scenes. The best results are shown in bold in each group.

### v2

@@ 2.2.3 Cluster-based Local Matching
After extracting the global geometric embedding, we implement a cluster-based matching module to obtain the initial intra-domain 2D-3D matches. This cluster-based GNN  [45] has been shown to be more computationally efficient than its complete-graph counterpart [38]. The network considers the local point features from both $\mathcal{\widetilde{P}}$ and $\mathcal{\widetilde{Q}}$ a complete set, then clusters the feature with strong correlations into the same group and restricts the message passing within each group. In addition to its low computational complexity, we found that cluster GNN can effectively utilize global-to-local geometric cues, as the clustering operation inherits the property of global graph clustering and forces it to distinguish ambiguous local features even with similar global embedding.
Methods ScanNet [10] MegaDepth [24]
BPnPNet [5] 10.00 / 10.00 / 10.02 199.17 / 128.90 / 154.68 4.35 / 6.82 / 9.86 13.60 10.22 / 10.63 / 10.89 16.13 / 32.01 / 61.58 1.67 / 3.17 / 5.44 12.95
GoMatch [56] 11.18 / 11.23 / 18.01 112.69 / 112.78 / 136.50 0.19 / 0.91 / 2.63 13.18 15.67 / 22.43 / 28.01 10.60 / 10.08 / 34.63 0.06 / 1.06 / 3.73 14.94
k=10 BPnPNet [5] 10.00 / 10.00 / 10.03 104.68 / 135.94 / 160.54 4.67 / 7.30 / 10.92 10.84 10.36 / 10.72 / 10.97 16.63 / 34.69 / 67.77 1.64 / 3.30 / 5.97 10.74
GoMatch [56] 10.91 / 18.98 / 31.12 111.18 / 114.94 / 128.97 10.08 / 0.35 / 2.08 14.25 18.90 / 35.67 / 44.99 10.18 / 11.29 / 16.65 0.02 / 0.12 / 1.92 18.76
Table 1: 2D-3D Matching. We present AUC scores for reprojection errors threshold at 1, 5, and 10 pixels; rotation and translation error quantiles at 25, 50, and 75%; and matching precision. Parameter $k$ is the number of images retrieved from the database to narrow down the search space. The best results are bold. DGC-GNN nearly doubles the AUC scores of GoMatch and reduces the pose errors to their ${\approx}33\%$.
@@
Initialization. As an initialization procedure for the cluster attention module, we run the general self and cross-attention modules proposed in GoMatch [56]. For each local point $\mathbf{b}_{\mathbf{p},n}$, we construct a local graph according to its $k^{\prime}$ nearest neighbours in the Euclidean space and update the associated feature $\mathbf{\widetilde{f}}_{\mathbf{p},n}\in\mathbb{R}^{2d}$ by Eq. 5. Note that we ignore the angular embedding at this stage due to the unaffordable memory requirements with space complexity $\mathcal{O}(Nk^{\prime 2})$, where $N$ is the number of local points. We then use linear attention [21, 49] as a cross-attention mechanism, which allows each point in one modality to interact with all points from another modality. This not only facilitates inter-modality in the feature matching but also reduces the computational complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$.
@@
Cluster-based Attention. After the graph initialization, the features $\mathbf{\widetilde{f}}_{\mathbf{p}}$ and $\mathbf{\widetilde{f}}_{\mathbf{q}}$ coming from the image and the point cloud respectively, are concatenated and processed in a two-level hierarchical clustering attention module. The hierarchical structure is effective in suppressing erroneous groupings. At the first level, we cluster the feature vectors into $I$ coarse groups. In the second level, each coarse group is divided into several small groups. The local point information exchange is conducted at each level and only within each group to obtain more representative features. After the sparse clustering, each feature vector is transformed back to its original position and then split again into $\mathbf{\widetilde{f}^{\prime}}_{\mathbf{p}}$ and $\mathbf{\widetilde{f}^{\prime}}_{\mathbf{q}}$ to obtain the keypoints both in the 2D and 3D spaces.
@@
$\mathcal{M}_{init}=\{(\widetilde{n},\widetilde{m})\ |\ \forall(\widetilde{n},\widetilde{m})\in\text{MNN}(\mathcal{S})\},$ (8)
where MNN is the mutual nearest neighbors operator. Set $\mathcal{M}_{init}$ provides initial 2D-3D matches that we further filter in Sec. 2.3 to keep the accurate correspondences only.
@@ 2.3 Outlier Rejection
After obtaining the initial matches, outlier pruning runs to remove the incorrect ones. We apply the same outlier rejection network as in GoMatch [56], whose input is the concatenated 2D and 3D keypoint features $\mathbf{\widetilde{f}^{\prime}}_{\widetilde{n},\widetilde{m}}=\mathbf{\widetilde{f}^{\prime}}_{\mathbf{p},\widetilde{n}}\oplus\mathbf{\widetilde{f}^{\prime}}_{\mathbf{q},\widetilde{m}}$ and outputs the matching confidence of each matched pair. The final predicted matches are obtained as follows:
$\mathcal{M}_{final}=\{(\widetilde{n}^{\prime},\widetilde{m}^{\prime})\ |\forall\ \text{cls}(\mathbf{\widetilde{f}^{\prime}}_{\widetilde{n},\widetilde{m}}\;|\;(\widetilde{n},\widetilde{m})\in\mathcal{M}_{init})\geq\theta\},$ (9)
@@ 2.4 Training Loss
$\begin{split}\mathcal{L}_{ot}=-\frac{1}{|\mathcal{M}_{gt}|+|\mathcal{I}|+|\mathcal{J}|}(\sum\limits_{(n,m)\in\mathcal{M}_{gt}}\log\mathcal{\bar{S}}_{n,m}+\\ \sum\limits_{i\in\mathcal{I}}\log\mathcal{\bar{S}}_{i,m+1}+\sum\limits_{j\in\mathcal{J}}\log\mathcal{\bar{S}}_{N+1,j}).\end{split}$ (10)
@@
where $w_{i}$ is the weight balancing the positive and negative samples, $y_{i}$ is the ground truth matching label for the $i$-th correspondences, $p_{i}$ is the predicted probability of a true match for the $i$-th correspondences. The total loss is the sum of the two terms as $\mathcal{L}=\mathcal{L}_{ot}+\mathcal{L}_{or}$.
@@ 3 Experiments
Training. We train the indoor model of DGC-GNN on the ScanNet [10] dataset and the outdoor model on the MegaDepth [24] dataset. We extract up to 1024 keypoints for each training image by the SIFT detector [27]. Similarly as in GoMatch, we first select a subset of the point cloud by applying image retrieval approaches [1, 50] to obtain potential images observing the same part of the scene as the input one. We randomly sample the retrieval pairs with a visual overlap of more than 35% on MegaDepth and 65% on ScanNet to ensure enough matches on each pair. For the global geometric embedding, we cluster the 2D/3D bearing vectors into $X=10$ groups, and each cluster center is connected to its $k=4$ nearest neighbors to build the global graph. For the local point graph, we connect each point with its 10 nearest neighbors and the cluster-based attentions are performed twice to force the intra-cluster information exchange.
@@
We use Adam optimizer with a learning rate of 1e-3. We train DGC-GNN with one 32GB Telsa V100 GPU. The convergence of the model typically requires 50 epochs.
@@
Datasets. We use ScanNet and MegaDepth for training and 2D-3D matching task evaluation. As a downstream application, we perform visual localization on the 7Scenes [46] and Cambridge Landmarks [22] datasets. MegaDepth is a popular outdoor dataset with 196 scenes captured around the world. The sparse 3D reconstructions are provided by the COLMAP [43] structure-from-motion software. Following [56], we train our outdoor model on 99 scenes and evaluate it with 53 scenes. ScanNet is a large-scale RGB-D indoor dataset comprising 1613 scans with over 2.5 million images. We randomly selected 105 scenes for the training and 30 for the evaluation. Cambridge Landmarks is a middle-scale outdoor dataset consisting of 6 individual scenes. A structure-from-motion algorithm provides the ground truth camera poses. We follow [22, 56] to evaluate our method on four scenes. 7Scenes is a small indoor dataset with RGB-D images and camera poses provided by the depth SLAM system. We evaluate on the standard test sequences.
@@
Evaluation Protocol. For matching on ScanNet and MegaDepth, we follow [56] and report the AUC score calculated from the reprojection errors. To calculate the errors for the 2D-3D matches in $\mathcal{M}_{final}$, we project the 3D points to the image plane using the ground truth and estimated camera poses. Then, we calculate the $L_{2}$ distance of the ground truth and estimated reprojected 2D points. We use multiple thresholds, 1, 5, and 10 pixels, to evaluate the AUC scores. The camera translation and rotation error quantiles at 25%, 50%, and 75% are also reported. Moreover, we evaluate the matching quality by calculating the matching precision P, which is the ratio of inlier matches after PnP-RANSAC to the number of final matches $\mathcal{M}_{final}$. For visual localization tasks, we report the median translation (in meters) and rotation (in degrees) camera pose errors.
@@
We compare with the two descriptor-free matchers GoMatch [56] and BPnPnet [5]. At inference, we use the 3D points from the top-$k$ retrieved database images to match with the keypoints from query images. Following [56], we report the upper bound of the AUC score using the ground truth matches. We refer to these values as Oracle. We select the GT matches by thresholding the reprojection error based on normalized image coordinates, using a threshold of $0.001$, to bypass the influence of camera intrinsics during GT selection. This is in contrast to what is done in [56]. Results on GT selected by a pixel threshold are in the supp. material. We use the official code with the default setting to generate the evaluation dataset on MegaDepth [24] and rerun GoMatch and BPnPNet with the released models. Note that we also tested GoMatch after retraining it on MegaDepth and achieved similar results as with the released model.
@@
Table 2: Ablation Study. AUC scores thresholded at 1, 5, and 10 pixels; rotation and translation error quantiles at 25, 50, 75% with the proposed components added one by one to the GoMatch pipeline on the MegaDepth dataset.
Matching Results. The results with $k=1$ and $k=10$ are presented in Table 1. Parameter $k$ is the number of retrieved image pairs that are used for evaluation. The proposed method outperforms GoMatch and BPnPNet by a significant margin on both scenes. Specifically, DGC-GNN achieves 10.2 / 37.64 / 44.04% reprojection AUC compared to GoMatch with 5.67 / 22.43 / 28.01% on MegaDepth with $k=1$. DGC-GNN halves the rotation and translation errors of GoMatch on all thresholds and it obtains better matching quality. Notably, the performance of DGC-GNN with $k=1$ surpasses that of GoMatch with $k=10$, indicating the effectiveness of our method even with a single view.
Figure 4: Outlier Sensitivity. The AUC scores of BPnPNet [5], GoMatch [56], and the proposed DGC-GNN thresholded at 1, 5, and 10 pixels are plotted as a function of the outlier ratio. Oracle represents the AUC upper bound using ground truth matches.
Sensitivity to Outliers. To evaluate the sensitivity to keypoint outliers, we follow the procedure in GoMatch [56]. The outliers are controlled by the outlier ratio, ranging from 0 to 1, calculated as the number of unmatched keypoints divided by the maximum of the numbers of 2D and 3D points. If the outlier ratio is 00, all the input 2D and 3D points are selected from the ground truth matches, and no outliers are included in the matching process. When it is $1$, we directly use the keypoints from the query image and 3D points from the top-$k$ retrieved images without any filtering or outlier removal. The results are shown in Fig. 4. Even in the presence of outliers, DGC-GNN outperforms other methods by a large margin. This indicates that our method is more robust to outliers and can handle challenging matching scenarios more effectively than the state-of-the-art.
Ablation Study. We investigate the effectiveness of different components of DGC-GNN on the 2D-3D matching quality on the MegaDepth dataset [24] with $k=10$. The results are reported in Table 2. We provide results with $k=1$ in the supp. material. We conduct the ablations by gradually adding the components: global geometric embedding (G. Emb), cluster attention (C. Att.), Color, and Angular embedding (Ang.) to the original GoMatch pipeline. Incorporating color information into the matching process significantly impacts the performance, resulting in improvements of 2.55 / 3.88 / 2.55% (AUC@1 / 5 / 10px). This demonstrates the importance of color cues for accurate and robust matching. The global-to-local geometric (G. Emb.) and Angular relation embedding (Ang.) substantially improve the matching performance by 1.90 / 5.51 / 5.52% and 1.11 / 3.36 / 3.47%, respectively. It highlights the effectiveness of incorporating global geometric context and local geometric details. The cluster attention mechanism also plays a vital role, improving performance by 0.78 / 3.28 / 3.48%. The best results are obtained when all components are added to the pipeline.
Methods No Desc. Maint. Privacy Cambridge-Landmarks [22] (cm, ${}^{\circ}$) MB used 7Scenes [46] (cm, ${}^{\circ}$) MB used
@@
E2E MS-Trans. [44] 83 / 1.47 181 / 2.39 86 / 3.07 162 / 3.99 1171 11 / 4.66 24 / 9.60 14 / 12.19 17 / 5.66 18 / 4.44 17 / 5.94 26 / 8.45 11171
DSAC* [3] 15 / 0.30 121 / 0.40 15 / 0.30 113 / 0.40 1112 12 / 1.10 12 / 1.24 11 / 1.82 13 / 1.15 14 / 1.34 14 / 1.68 13 / 1.16 11196
HSCNet [23] 18 / 0.30 119 / 0.30 16 / 0.30 119 / 0.30 1592 12 / 0.70 12 / 0.90 11 / 0.90 13 / 0.80 14 / 1.00 14 / 1.20 13 / 0.80 11036
DB HybridSC [6] 81 / 0.59 175 / 1.01 19 / 0.54 150 / 0.49 1113 - - - - - - - -
AS [41] 13 / 0.22 120 / 0.36 14 / 0.21 118 / 0.25 1813 13 / 0.87 12 / 1.01 11 / 0.82 14 / 1.15 17 / 1.69 15 / 1.72 4 / 1.01 -
SP [11]+SG [38] 12 / 0.20 115 / 0.30 14 / 0.20 117 / 0.21 3215 12 / 0.85 12 / 0.94 11 / 0.75 13 / 0.92 15 / 1.30 14 / 1.40 15 / 1.47 22977
DF GoMatch [56] 25 / 0.64 283 / 8.14 48 / 4.77 335 / 9.94 1148 14 / 1.65 13 / 3.86 19 / 5.17 11 / 2.48 16 / 3.32 13 / 2.84 89 / 21.12 11302
@@
Table 3: Visual Localization. We report the median pose errors (cm, ${}^{\circ}$) and storage requirements (MB) on the scenes of the 7Scenes [46] and Cambridge-Landmarks [22] datasets. Three groups of methods are shown: end-to-end (E2E), descriptor-based (DB), and descriptor-free (DF). We do not show BPnPNet as it fails on most scenes. The best results are shown in bold in each group.